# Enumerative Geometry of Quantum Periods

### v1

@@ A.2. Bell polynomial identities
For $a_{0}=0$ the claimed equation reduces to $[x^{m}]f^{n}=\widehat{B}_{m,n}(a)$, which is Proposition A.8. In general, we can write $f(x)=a_{0}+g(x)$, where $g(x)=\sum_{k>0}a_{k}x^{k}$ has no constant term. Then
@@
(A.1) $\widehat{B}_{n,k}(a_{1},\ldots,a_{n-k+1})=\frac{k!}{n!}B_{n,k}(1!a_{1},\ldots,(n-k+1)!a_{n-k+1}).$
@@
We substitute $x_{i}=i!a_{i}$ and use (A.1) to find that $\binom{\tau}{k}B_{n,k}(a)$ equals
@@
By Lemma A.9, the right hand side of the claim can be written as
@@
In the last equality we used Lemma A.9 and in the one before we used Lemma A.10 for the parameters $n-1,k,\tau=n$ and $\alpha(l,m)=m+1$. ∎
Proposition A.12.
@@
that we recognize as the statement of Lemma A.11 which concludes the proof. ∎

### v2

@@ B.2. Bell polynomial identities
For $a_{0}=0$ the claimed equation reduces to $[x^{m}]f^{n}=\widehat{B}_{m,n}(a)$, which is Proposition B.8. In general, we can write $f(x)=a_{0}+g(x)$, where $g(x)=\sum_{k>0}a_{k}x^{k}$ has no constant term. Then
@@
(B.1) $\widehat{B}_{n,k}(a_{1},\ldots,a_{n-k+1})=\frac{k!}{n!}B_{n,k}(1!a_{1},\ldots,(n-k+1)!a_{n-k+1}).$
@@
We substitute $x_{i}=i!a_{i}$ and use (B.1) to find that $\binom{\tau}{k}B_{n,k}(a)$ equals
@@
By Lemma B.9, the right hand side of the claim can be written as
@@
In the last equality we used Lemma B.9 and in the one before we used Lemma B.10 for the parameters $n-1,k,\tau=n$ and $\alpha(l,m)=m+1$. ∎
Proposition B.12.
@@
that we recognize as the statement of Lemma B.11 which concludes the proof. ∎